### v3

Abstract
When learning stable linear dynamical systems from data, three important properties are desirable: i) predictive accuracy, ii) provable stability, and iii) computational efficiency. Unconstrained minimization of reconstruction errors leads to high accuracy and efficiency but cannot guarantee stability. Existing methods to remedy this focus on enforcing stability while also ensuring accuracy, but do so only at the cost of increased computation. In this work, we investigate if a straightforward approach can simultaneously offer all three desiderata of learning stable linear systems. Specifically, we consider a post-hoc approach that manipulates the spectrum of the learned system matrix after it is learned in an unconstrained fashion. We call this approach spectrum clipping (SC) as it involves eigen decomposition and subsequent reconstruction of the system matrix after clipping all of its eigenvalues that are larger than one to one (without altering the eigenvectors). Through detailed experiments involving two different applications and publicly available benchmark datasets, we demonstrate that this simple technique can simultaneously learn highly accurate linear systems that are provably stable. Notably, we demonstrate that SC can achieve similar or better performance than strong baselines while being orders-of-magnitude faster. We also show that SC can be readily combined with Koopman operators to learn stable nonlinear dynamics, such as those underlying complex dexterous manipulation skills involving multi-fingered robotic hands. Further, we find that SC can learn stable robot policies even when the training data includes unsuccessful or truncated demonstrations. Our codes and dataset can be found at https://github.com/GT-STAR-Lab/spec_clip.
keywords:
Linear Dynamical Systems, Stability, System Identification.
1 Introduction
In spite of the growing dominance of deep neural networks in various fields, modeling and learning linear dynamical systems remain relevant due to the availability for analytical solutions, significantly lower computational burden, and ease of analyses and inspection. They continue to find use in diverse applications such as computer vision (Chan and Vasconcelos, 2005; Boots et al., 2007) and time series prediction (Liu et al., 2024; Takeishi et al., 2017). Even when the underlying dynamics are nonlinear, researchers are increasingly utilizing tools from linear systems theory by leveraging the Koopman Operator theory (Koopman, 1931; Williams et al., 2015) to approximately represent nonlinear systems as linear systems in a higher-dimensional lifted state space. These techniques have further increased the relevance and importance of learning linear systems in complex applications, such as differential drive robots (Abraham et al., 2017), serial-link manipulators (Abraham and Murphey, 2019), legged robots (Kim et al., 2024; Li et al., 2024), soft robotic manipulators (Bruder et al., 2021), and dexterous robotic manipulation (Han et al., 2023, 2024; Chen et al., 2024).
When time-series data associated with system evolution is available, learning such linear systems can be considered as a form of self-supervised learning and can be solved efficiently by minimizing a regression objective (Ljung, ). However, such unconstrained optimization methods tend to overlook the important properties of the underlying linear systems, such as stability. This is particularly critical in linear system modeling, as a linear system comprises only three evolution modes: convergence, divergence, and oscillation (Robinson, 2012). Consequently, even if the underlying system dynamics are stable, the stability of learned system could be highly sensitive to the training data distribution, potentially leading to unstable solutions and catastrophic long-term predictions.
A number of methods have been developed to enforce stability of learned linear systems and reduce sensitivity to training data (Lacy and Bernstein, 2002, 2003; Boots et al., 2007; Huang et al., 2016; Gillis et al., 2020; Mamakoukas et al., 2020, 2023). These methods take one of three main approaches: a) minimizing the regression objective under the stability constraints (Lacy and Bernstein, 2002, 2003), b) iterating between solving an unconstrained problem and optimizing its solution under the stability constraints (Boots et al., 2007; Huang et al., 2016), and c). iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective (Gillis et al., 2020; Mamakoukas et al., 2020, 2023). Regardless of the underlying approach, these methods inevitably increase the computation burden, potentially counteracting one of the primary advantages of leveraging linear techniques. Furthermore, they exhibit poor scalability when applied to high-dimensional systems, which are necessary in both modeling inherently high-dimensional state spaces (e.g., images) and effectively approximating complex nonlinear systems using Koopman operators Koopman (1931) (e.g., robotic systems).
In this work, we suggest and investigate a simple yet surprisingly effective method, Spectrum Clipping (SC), to learn the stable linear systems from data requiring negligible additional computations. This approach involves a straightforward implementation: 1) compute the system evolution matrix by solving the least-squares regression problem, 2) perform eigen decomposition of the system matrix and extract all eigenvalues, 3) clip eigenvalues with magnitudes larger than one to one, and 4) reconstruct the system evolution matrix using the clipped eigenvalues and the unchanged eigenvectors. SC enforces stability by construction but neither requires complex constrained optimization nor updating the matrix iteratively. As a result, SC is both efficient and scales well to high-dimensional systems. We are inspired to investigate this simple approach since strategies related to clipping a matrix’s spectrum have proven effective in other domains. For instance, directly controlling the singular values of linear layers seems to enhance training stability and reduce the model’s generalization error (Boroojeny et al., 2024; Sedghi et al., 2018; Senderovich et al., 2022). We are also inspired by seminal works that revealed the spectral properties of the Koopman operator (Rowley et al., 2009; Mezić, 2013; Mezic, 2020). Indeed, if one clips the spectrum of a learned Koopman matrix, this simple approach can help enforce stability when learning nonlinear systems.
We designed and carried out thorough experiments involving diverse applications and benchmark datasets to investigate the effectiveness of SC and compared its performance against strong baselines. We specifically compared SC with three baselines: CG (Boots et al., 2007), WLS (Huang et al., 2016), and SOC (Mamakoukas et al., 2020). We first evaluated SC’s ability to learn stable autonomous linear systems by considering the challenge of learning dynamic textures from videos from two benchmark datasets (Chan and Vasconcelos, 2005; Hadji and Wildes, 2018). We also evaluated SC’s ability to learn systems with control inputs by evaluating its performance when learning the dynamic model of a Franka Emika Panda robotic arm from a standardized dataset (Gaz et al., 2019). Our results across all experiments demonstrate that SC achieves the lowest reconstruction error and generates stable long-term predictions. Most notably, SC is able to achieve this level of performance while being orders-of-magnitude faster in learning provably-stable linear systems.
We also investigated if SC can be combined with Koopman operators to learn complex nonlinear systems. Specifically, we evaluated SC’s ability to learn stable linear systems in the lifted space that encode complex dexterous manipulation skills for a multi-fingered robotic hand (Han et al., 2023). Our analyses reveals that unconstrained learning produces unstable linear systems if the training data includes unsuccessful demonstrations or when the task horizon is truncated. We are motivated by the fact that such characteristics are commonly observed when robot learning data is collected from humans (Qin et al., 2022; Shaw et al., 2023; Bahl et al., 2022). Our results show that SC can be readily combined with Koopman operators to learn stable nonlinear systems in mere seconds, while remaining robust to the undesirable characteristics in the training data. We also find that unstable and erratic robot behavior can arise due to a subset of the underlying dynamic modes (see Schmid (2022)). We find that clipping the spectrum of the learned Koopman matrix effectively stabilizes its rollouts and enables the robot to safely complete the task. These findings are in line with the well-established connections between Koopman modes and the state-space geometry of the underlying dynamical systems (Rowley et al., 2009; Mezić, 2013; Mezic, 2020).
2 Problem Formulation
Consider a discrete-time autonomous linear system as
$\mathrm{x}_{t+1}=A\mathrm{x}_{t},$ (1)
where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the system state at time $t$, and $A\in\mathcal{S}_{A}\subset\mathbb{R}^{n\times n}$ is the system evolution matrix. Here, we use $\mathcal{S}_{A}$ to denote the solution space of the matrices $A$.
Further, suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T}]$. Let $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, then to learn the linear system from the dataset $D$ is to minimize the regression objective as follow:
$A=\arg\min_{\mathcal{S}_{A}}||Y-AX||_{F}^{2},$ (2)
Where $||\cdot||_{F}$ is the matrix Frobenius norm. Now, optimizing the $A$ matrices amounts to solving a least-square regression problem (Montgomery et al., 2021; Seber and Lee, 2012).
However, optimizing the Equation. 2 does not impose any constraints on the matrix $A$, such as the system stability that is crucial for the long-term prediction. Long-term prediction is particularly important for various applications, such as ensuring safety in robotic manipulation (Han et al., 2023, 2024). Therefore, we need to consider the stability guarantee constraint (Boots et al., 2007; Huang et al., 2016). Let $\{\lambda_{i}(A)\}^{n}_{i=1}$ be the eigenvalues of the matrix $A$, such matrix $A$ is stable if it has $\max\{\lambda_{i}(A)\}^{n}_{i=1}<=1$ (i.e., all the eigenvalues have a magnitude of one at most). Now, we introduce $\mathbb{S}_{A}=\{A\in\mathcal{S}_{A}|\max\{\lambda_{i}(A)\}^{n}_{i=1}<=1\}$ to denote the solution space where matrix $A$ is stable.
Learning Stable Linear Systems from Data:
Our goal is to learn the stable linear systems from the dataset $D$. Therefore, the objective in Equation. 2 turns into
$A=\arg\min_{\mathbb{S}_{A}}||Y-AX||_{F}^{2},$ (3)
which is highly non-convex (Boyd and Vandenberghe, 2004; Mamakoukas et al., 2020). In practice, obtaining an analytical solution to Equation 3 is impractical and necessitates approximate methods.
3 Spectrum Clipping for Stability
In this paper, we evaluate a simple method, which we refer to as Spectrum Clipping (SC), that effectively learns stable linear systems while being significantly more computational efficient.
In this method, we first compute the system evolution matrix $A_{ls}$ by optimizing the least square regression (Eqn. 2). Then, we perform eigen decomposition of the matrix $A_{ls}$
$A_{ls}=M\times\Lambda\times M^{-1}$ (4)
where $\Lambda$ is a diagonal matrix with all the eigenvalues of $A_{ls}$ on its diagonal, and $M$ is the matrix of corresponding eigenvectors (modal matrix).
There might be some defective eigenvalues whose algebraic multiplicity is larger than its geometric multiplicity. In this case, we may assign linear dependent eigenvectors to them but still keep the $\Lambda$ matrix diagonal. Although this may result in a non-invertible modal matrix as well as a slight numerical difference between the original matrix $A_{ls}$ and the reconstructed matrix $\bar{A_{ls}}=M\times\Lambda\times M^{-1}$, the deviation is negligible w.r.t the task performance in our experiments (an exception is discussed in section. 5).
For those unstable eigenvalues (i.e., eigenvalues whose norm is larger than 1) in matrix $\Lambda$, we clip their values in one of the following two ways.
Spectrum Clipping to 1
we clip them to exactly $1$ and constitute a new matrix $\bar{\Lambda}$. Finally, we multiply $\bar{\Lambda}$ with the original eigenvector matrices and get a modified matrix $\bar{A_{sc}}=M\times\bar{\Lambda}\times M^{-1},\bar{A_{sc}}\in\mathbb{S}_{A}$. This strategy allows us to obtain a linearly stable matrix that maintains maximum expressivity, which is also beneficial for long-term prediction. For instance, when learning dynamic textures, expressiveness is essential for accurate prediction beyond the training data horizon.
Spectrum Clipping to 1 - $\bm{\varepsilon}$, $\bm{\varepsilon\in}$ (0, 1]
By selecting $\varepsilon>0$, e.g., $\varepsilon=10^{-5}$, the resulting system stability meets the criteria of well-known Lyapunov stability [^1] All eigenvalues have magnitudes strictly less than one.. However, this choice might limit the system’s expressivity, and detailed discussions are included in the Appendix. D. Similarly, we then recompose the clipped $\bar{\Lambda}$ matrix with modal matrix $M$ and obtain $\bar{A_{sc}}$.
4 Empirical Analyses
In this section, we evaluate SC in terms of computation efficiency, stability, and expressivity when learning LDS from data and compare its performance against existing methods. Note that for all experiments, we clip the spectrum to one (see Appendix. D for the effects of different $\varepsilon$ values).
4.1 Experimental Design
Computation Platform:
We conduct all experiments on Alienware Aurora R13, with 10-core 12th Gen Intel Core i7-12700KF 1.0-GHz CPU and a 32G RAM.
Baselines:
We compare the performance of SC against four baselines: i) LS: vanilla least squares, ii) (CG): constraint generation (Boots et al., 2007),iii) WLS: weighted least squares (Huang et al., 2016), and iv) SOC: a characterization approach (Mamakoukas et al., 2020, 2023).
Datasets:
We evaluate all approaches on the following datasets.
USCD (Chan and Vasconcelos, 2005): A dataset of 254 highway traffic videos. Each video has 48-52 frames and the size of each frame is $48\times 48$. To lower the dimension, we apply the SVD method on each sequence to obtain a subspace dimensions $r\in\{3,30\}$, as done by Mamakoukas et al. (2020).
DTDB (Hadji and Wildes, 2018): A dataset of 285 dynamic texture movements, such as turbulence, wavy motions, etc. Each video consists of varying frames, and we also apply the SVD method to lower the state dimension to $r=300$.
Note that due to the high dimension of the state space in the DTDB datasets, only LS, SOC, and SC are applicable. Both CG and WLS involve solving a Kronecker product, which has a space complexity of $O(d^{4})$ ($d$ is the state space dimension), leading to memory overflow.
Metrics:
We evaluate all approaches in terms of: i) Computation time: Time taken to train a stable LDS, ii) Memory Usage: Memory used during the entire computation process, and iii) Reconstruction error: The $L1$ distance between generated state sequences and ground truth states.
4.2 SC significantly reduces computation & memory burden
In this subsection, we compare the training efficiency among each approach. We first report the average computation time across the 254 input sequences of the UCSD dataset at each dimension in Fig. 1 (a). The results show that SC is significantly more time-efficient than WLS, CG, and SOC. This is because that SC only needs the eigen-decomposition and eigen-recomposition to analytically compute the stable LDSs, while others require to iteratively find a solution. Note that though LS is the most time-efficient, it can not provide with stability.
In Fig. 1 (b), we then report the results on the DTDB dataset. The results also indicate that the computation time of SC is significantly and consistently shorter than that of SOC, showing that SC is more scalable w.r.t. the input dimensions. Similarly, the memory usage results reported in Fig. 1 (c) and Fig .1 (d) suggest that SC requires significantly less memory compared to CG, WLS, and SOC.
Figure 1: (a): Computation time on UCSD dataset (solid lines indicate the median); (b): The boxplot of the average computation time on DTDB dataset; (c): Memory usage on UCSD dataset (solid lines indicate the median); (d): The boxplot of the average memory usage on DTDB dataset.
4.3 SC effectively identifies the underlying LDSs
In this subsection, we compare the prediction accuracy of the identified LDSs by each approach. We first present the average reconstruction error across all states at each time step for the UCSD dataset in Fig. 3. For illustrative purposes, we select $r=3,20$ and $30$ as examples. It is clearly shown that as the input dimension increases (e.g., $r=30$), LS fails to reconstruct the input states. However, SC consistently achieves results comparable to CG, WLS, and SOC, while being orders of magnitude faster for computation and highly efficient in memory usage (see Fig. 1 (a) and Fig. 1 (c)).
We then show the average reconstruction error for the DTDB dataset in Fig. 3. Given the varying time steps of the input sequences, the reconstruction error is averaged over the time percentiles. The results indicate that SC outperforms SOC since it achieves a lower reconstruction error. Note that though the LS approach achieves the lowest error within the training data horizon, it does not guarantee stability, which is crucial for long-term prediction (see Section. 4.4).
Figure 2: Average reconstruction error evaluated on different subspace dimensions of the UCSD dataset.
Figure 3: Average reconstruction error evaluated at each time step percentile of the DTDB dataset.
4.4 SC enables stable long-term predictions
In this subsection, we demonstrate that the SC approach learns LDSs capable of making long-term dynamic predictions that extend beyond the training data horizon. To verify this, we extend the rollout horizon to $T=3000$ time steps for the 285 LDSs learned from the DTDB dataset. In this case, an accurate LDS should be able to repeat the dynamical pattern learned from the training data without either blowing up or converging during the extended rollout horizon.
We first select two example sequences (i.e., flushing toilet and running stream) and show the qualitative results in Fig. 6. These results indicate that LS approach mostly learns the LDSs that blow up during the extended horizon, as it does not provide with stability guarantee. Conversely, the SOC approach tends to learn LDSs that converge to zeros, indicating a lack of dynamic behavior.
Figure 4: DTDB Sequence 185: A flushing toilet.
Figure 5: DTDB Sequence 205: A running stream.
Figure 6: Qualitative results for 2 example sequences from the DTDB dataset.
We then define several quantitative evaluation metrics, which are divided into two categories.
Magnitude: i) Turning black: if the pixels of the final frame are all zero, ii) Blowing up: if the pixels of the final frame blow up, and iii) Normal: neither Turning black nor Blowing up. These characterizations aim to determine if the system evolution of the LDSs is still meaningful.
Dynamics: i) Moving: if the last few frames still exhibit the dynamical pattern. This characterization aims to determine if the system evolution of the LDSs still reflects the dynamic pattern.
Turning black ($\downarrow$) Blowing up ($\downarrow$) Normal ($\uparrow$) Moving ($\uparrow$)
LS 6.67% 45.96% 47.37% 62.11%
SOC 23.16% 0% 76.84% 22.11%
SC 14.39% 0% 85.61% 54.39%
Table 1: Quantitative metrics (ratios) characterizing predictions from each method.
In Table. 1, we report the quantitative ratios in terms of the defined metrics over the 285 sequences in the DTDB dataset. Note that the values for LS method are marked in gray as this method does not guarantee stability. The results indicate that both SOC and SC learn stable LDSs without blow-ups thanks to the stability guarantee, but SOC is less effective at modeling long-term dynamic patterns, as the learned LDSs often converge to equilibrium states (i.e., not moving) or even complete zeros (i.e., turning black). On the other hand, although the LDSs learned by the LS approach exhibit the most dynamic patterns, they tend to be unstable, leading to blow-ups.
4.5 SC’s benefits extend to nonlinear systems when combined with the Koopman Operator
In this subsection, we demonstrate that when combined with the Koopman Operator theory (Koopman, 1931), SC can also be effectively and efficiently applied to highly nonlinear robot manipulation tasks. For this, we introduce a highly nonlinear robot manipulation dataset.
DexManip (Han et al., 2023; Rajeswaran et al., 2018b): A dataset of Adroit Hand performing two dexterous manipulation tasks in MUJOCO simulation environment (Todorov et al., 2012). Though the original state evolution is nonlinear, but as shown in Han et al. (2023), via lifted to a higher-dimension Hilbert space (e.g., $r\approx 750$), the system can be considered linear in the lifted state space. Two tasks are briefly described as follows: i) Tool Use: The robot hand is tasked with picking up the hammer and driving the nail into the board placed at a random height, and ii) Object Relocation: The robot hand is tasked with moving a cylinder to a random target location (see Appendix. A for the details of each dataset). Similar to the DTDB dataset, only LS, SOC, and SC are applicable due to the high dimensionality of the lifting space.
Computation Time (sec) Memory Usage (MB) Task Success Rate (%) Safety Rate (%)
Datasets LS SOC SC LS SOC SC LS SOC SC LS SOC SC
Tool 0.6 192.8 0.8 224.6 331.7 257.2 55.5 0.0 94.0 0.0 100.0 100.0
Reloc 0.7 6608.1 0.8 215.3 321.3 247.5 96.0 48.6 93.8 100.0 100.0 100.0
Table 2: SC not only alleviates the computational and memory burden on the DexManip datasets, but also enables successful and safe robot manipulation.
We first report the computation time and memory usage in Table. 2. As seen, SC also alleviates the computational and memory burden compared to the SOC method. Note that the values for LS method are marked in gray as this method does not guarantee stability. We then show the average reconstruction error on the DexManip datasets in Fig. 7 (a) & (b), where SC demonstrates the best performance, while LS results in significantly larger errors on the Tool dataset due to its lack of stability.
Figure 7: (a): Average reconstruction error on Tool dataset; (b): Average reconstruction error on Reloc dataset; (c): Average reconstruction error on Tool_short dataset; (d): Average reconstruction error on Reloc_with_failures dataset. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
To further underscore the importance of ensuring stability when learning LDSs for robotic dexterous manipulation tasks in the DexManip datasets, we also compute the task success rates and the safety rates (both reported in Table. 2) for each task when tracking the generated robot hand trajectories. The definition of these two metrics are included in Appendix. A.
For both tasks, SC learns stable LDSs that achieve the highest task success rates and complete the tasks safely. In contrast, LS fails to complete tasks safely, while SOC underperforms due to the fact it learns overly conservative LDSs that lack the ability to generate diverse trajectories.
4.6 SC can learn stable LDSs from shorter and corrupted datasets
In real-world settings, the collected demonstrations may be imperfect, unlike those in the original DexManip datasets. Hence, we introduce two additional datasets: (i) Tool_short, where each demonstration is shorter, posing challenges in learning stable LDSs with a limited horizon, and (ii) Reloc_with_failures, which includes a few failed demonstrations (see Appendix. A for more details).
Computation Time (sec) Memory Usage (MB) Task Success Rate (%) Safety Rate (%)
Datasets LS SOC SC LS SOC SC LS SOC SC LS SOC SC
Tool_short 0.3 3470.2 0.4 80.2 280.5 112.8 3.3 0.0 90.7 0.0 100.0 100.0
Reloc_with_failures 0.6 4564.6 0.9 216.1 322.1 248.4 9.6 0.0 94.4 0.0 100.0 100.0
Table 3: SC can efficiently learn stable manipulation skills from shorter and corrupted datasets.
In Table. 3, we report the computational time and memory usage, and it is evident that SC is more efficient than SOC. In Fig. 7 (c) & (d), we show the average reconstruction error over the original uncorrupted datasets. From these results, we can observe that the LDSs learned by the LS approach tend to be highly unstable, as evidenced by the significantly increasing error over time. This indicates the importance of enforcing stability in learned LDSs to effectively handle corrupted demonstrations.
A similar conclusion can be drawn from the task success rates and safety rates (both reported in Table. 3), where SC demonstrates significantly superior performance.
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in Rowley et al. (2009), we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset.
Specifically, for each eigenvalue $\lambda_{i}\ (i=1,...,n)$ of system matrix $K$, we can find corresponding eigenvector $\mathbf{v_{i}}$ (i.e., the mode) and eigenvector $\mathbf{w_{i}}$ of adjoint matrix $K^{*}$, such that $K\mathbf{v_{i}}=\lambda_{i}\mathbf{v_{i}}\text{, and }K^{*}\mathbf{w_{i}}=\overline{\lambda_{i}}\mathbf{w_{i}}$. Then following (Rowley et al., 2009), we construct the corresponding eigenfunction $\varphi_{i}:\mathbb{R}^{n}\rightarrow C$, defined as $\varphi_{i}(\mathbf{z})=\langle\mathbf{z},\mathbf{w}_{i}\rangle$. By doing so, the rollout of eigenfunctions can be represented by $\varphi_{i}(K\mathbf{z})=\lambda_{i}\varphi_{i}(\mathbf{z})$, and the states can be reconstructed by $\mathbf{z_{k}}=\sum_{i=1}^{n}\lambda_{i}^{k}\varphi_{i}(\mathbf{z_{0}})\mathbf{v_{i}}$. Based on this formula, specific modes can be selected for rollouts.
In our experiments, we first computed the Koopman matrix $K_{f}$ from $Reloc\_with\_failures$ dataset. We then sorted the eigenvalues by their norms from largest to smallest and divided them into two subsets for rollouts: (i) Unstable Set: the first 52 Koopman modes with unstable eigenvalues, which have norms greater than 1, and (ii) Stable Set: the remaining 707 modes. We then visualized the robot hand motions of each subset via the eigenfunction rollouts in Fig. 8, respectively. It can be seen that the rollout of the Unstable Set before clipping clearly generate unstable hand motions (see upper line of the green box), which results in the low success rate and safety rate when combined with Stable Set for the final task performance (as shown in Table. 3). Further, we clipped the matrix $K_{f}$ and visualized the rollouts of the clipped Unstable Set in Fig. 8. It is obvious that after clipping, the robot hand motions become stable (see bottom line of the green box), enabling the composite system to safely complete the task (as also shown in Table 3).
Figure 8: Rollout with two subsets of the Koopman modes. The green box represents rollouts using Unstable Set (norm of eigenvalue larger than 1). The blue box represents rollout with the rest Stable Set. In the green box, the upper line shows the rollout before clipping and the lower line shows it after clipping. It is obvious that after clipping, the rollout becomes stable, thereby allowing the composite rollout of both sets to safely complete the task (as shown in Table 3).
4.7 SC can identify systems with control inputs
In this subsection, we demonstrate the comparable capability of SC in identifying stable LDSs with control inputs (see Appendix. B for details). To this end, we first introduce another dataset that includes the control commands.
Franka Panda (Gaz et al., 2019): A dataset of Franka Panda Robot Arm’s movements. The robot state space has 17 dimensions ($r=17$), including the coordinates of end effector ($\mathbb{R}^{3}$), joint angle values ($\mathbb{R}^{7}$), and joint angular velocities ($\mathbb{R}^{7}$), and the control input has $\mathbb{R}^{7}$ joint torques. This dataset consists of 3100 time steps of control inputs (randomly generated) and the corresponding robot states.
To examine how dataset size affects performance, we adjust the dataset size to $d=100,2000$ and $3000$ samples. For each size, we randomly select five subsets from the whole dataset, resulting in totally 15 subsets. We then train LDSs, i.e., the system matrix $A$ and the control matrix $B$, using different approaches on all 15 subsets.
We first evaluate the long-term prediction accuracy of each approach. For this, we first command the random control inputs in simulation with ground-truth system dynamics to generate test trajectories. We then rollout the trained $A$ and $B$ matrices with the same control inputs to compute the Reconstruction error for each approach. From the results shown in Fig. 9, we can see that SC stills perform comparably to all other approaches.
Additionally, we use the trained $A$ and $B$ matrices to construct an LQR controller, and use such controller to track a figure-8 pattern in simulation (similar in Mamakoukas et al. (2020)). How accurately the figure is tracked indicates how well the learned LDS aligns with the actual dynamical system with control inputs. For each approach, we repeat this process five times with different initial positions of the end effector. We show the qualitative tracking performances in Appendix. C.
Figure 9: Distance to correct end effector position calculated at each time step. Solid lines represent median values, and shaded areas indicate quartile values over five policy rollouts.
5 Limitations and Future Work
Though spectrum clipping demonstrates its efficacy on identifying the stable LDSs over various datasets, there are still opportunities for future study. First, the performance on the dataset with control inputs is not yet fully comparable to other state-of-the-art approaches, indicating potential for algorithmic improvements. Second, while we demonstrate the effectiveness of SC experimentally, a deeper analysis of the spectrum properties before and after clipping could provide valuable theoretical insights. Third, if the system matrix is not diagonalizable, spectrum clipping may not be applicable. For example, if $A_{ls}=[1,1,0;0,1,1;0,1,1]$, the decomposition and recomposition algorithm introduced in Section. 3 can not maintain the original matrix. However, our experiments showed that a tiny Gaussian noise $N(0,10^{-24})$ added to each element of this matrix helped maintaining its structure after de- and re-composition. Therefore, this may not be a concern for the applications considered in this work, as learned matrices typically do not have a noise-free structure.
6 Conclusion
In this paper, we demonstrate that Spectrum Clipping is an efficient data-driven method for identifying stable LDSs. Through extensive experiments across multiple datasets, we show that it is not only significantly more time- and memory-efficient but also achieve superior performance. Furthermore, we demonstrate that it also excels in long-term dynamics prediction. Additionally, it can be applied to the highly nonlinear robot manipulation tasks with the Koopman Operator theory and the linear dynamical systems with control inputs.
Appendices
Appendix A Details of DexManip tasks
DexManip is a dataset collected when using Adroit Hand to conduct two different dexterous manipulation tasks in MUJOCO simulation environment. In this dataset, the states of hand and states of object is captured, concatenated and lifted to a higher-dimension Hilbert space at each time step (see Han et al. (2023)), so that the hand-object system becomes an LDS in the Hilbert space. Thereafter, we use a simple PD controller to track the generated hand trajectories in the simulator. The tasks are detailed as follows.
Tool Use: This task is mainly about picking up the hammer to drive the nail into the board placed at some height. The hand and object states are lifted to a higher-dimensional $z$ states and different approaches are used to learn the LDSs in the $z$ state space. There are two types of datasets:
  • Tool: This dataset consists of 200 rollouts that continues after the nail is driven into the board, so that the hammer stays stable in the air ($T=100$). This is the original setting proposed in Rajeswaran et al. (2018a).
  • Tool_short: This dataset consists of 200 rollouts that end right after the nail is driven in ($T=35$).
This task is considered i) successful if at last time step (note that during evaluation, the policy will always run for 100 time steps, similar to the Tool dataset), the Euclidean distance between the final nail position and the goal nail position is smaller than 0.01m (Han et al., 2023), and ii) safe if the final distance from the hammer to the goal nail position is smaller than 0.3m, indicating that the robot hand neither drifts away nor throws away the hammer.
Object Relocation: This task is mainly about moving a cylinder to a randomized target location. Similarly, all approaches are operated in the higher-dimensional $z$ state space. The two types of datasets are:
  • Reloc: This dataset consists of 177 rollouts that objects are all moved to the target position successfully in the end ($T=100$).
  • Reloc_with_failures: In addition to the Reloc dataset, this one also includes the failure cases where the objects are failed to move to the target position in the end ($T=100$).
At each time step, if the Euclidean distance between the current cylinder position and the target position is smaller than 0.10m, then we have $\rho(t)=1$. This task is considered i) successful if $\sum_{t=1}^{T=100}\rho(t)>10$ (this criterion is the same as that used in Han et al. (2023)), and ii) safe if the robot hand is still near the target position (i.e., the distance is smaller than 0.5m) at final time step, indicating that the robot hand does not drifts away.
Appendix B Spectrum Clipping for Controlled System
Consider a discrete-time controlled linear system as
$\mathrm{x}_{t+1}=A\mathrm{x}_{t}+B\mathrm{u}_{t},$ (5)
where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n},\mathrm{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ are the system state and control signal at time $t$, respectively. Similarly, $\{A,B\}\in\mathcal{S}_{A,B}\subset\{\mathbb{R}^{n\times n},\mathbb{R}^{n\times m}\}$, where $A,B$ are the system evolution matrix and the control matrix. We use $\mathcal{S}_{A,B}$ to denote the solution space of the matrices $A$ and $B$ that describes the system evolution of Equation. 5.
Now, suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{u}_{1},\mathrm{x}_{2},\mathrm{u}_{2},\cdots,\mathrm{x}_{T}]$. Let $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, and $U=[\mathrm{u}_{1},\mathrm{u}_{2},\cdots,\mathrm{u}_{T-1}]\in\mathbb{R}^{m\times T-1}$, then to learn the linear system from the dataset $D$ is to minimize the regression objective as follow:
$A,B=\arg\min_{\mathcal{S}_{A,B}}||Y-AX-BU||_{F}^{2},$ (6)
Similarly, optimizing Equation. 6 does not guarantee stability. Here, the stability constraint for the controlled system is similar to that of an autonomous system, with the exception that the solution space changes to $\mathbb{S}_{A,B}=\{\{A,B\}\in\mathcal{S}_{A,B}|\max\{\lambda_{i}(A)\}^{n}_{i=1}<=1\}$. Therefore, the objective in Equation. 6 turns into
$A,B=\arg\min_{\mathbb{S}_{A,B}}||Y-AX-BU||_{F}^{2},$ (7)
Learning Stable Linear Systems from Data with Control Inputs:
To find the approximately optimal solution to Eqn. 7, to We first compute $A_{ls}$ and $B_{ls}$ by optimizing the least square regression (Eqn. 6). Then we clip matrix $A_{ls}$ to $\bar{A_{sc}}$ as described in Section. 3, while keeping $B_{ls}$ unchanged.
Appendix C Qualitative performance on the Franka Panda dataset
From the results shown in Fig. 10, we can observe that CG and SOC overall exhibit the best performance, especially when using only 100 training samples. However, as dataset size increases, the performance of SC is significantly improved, and become comparable to both CG and SOC, while still being orders of magnitude faster for training.
Figure 10: Rollouts of the end effector trajectory. Each subplot contains 25 trajectories, representing 5 different initial positions with matrices trained on 5 subsets.
Appendix D Tradeoff between stability and expressivity
As mentioned in Sec. 3, we can also select the $\varepsilon$ values to ensure that the resulting linear system meets the criteria for well-known Lyapunov stability. Through the experiments conducted on UCSD, DTDB, DexManip, and Franka Panda datasets, we demonstrate that this strategy does not only guarantee the Lyapunov stability, but also achieve comparable performance, provided that an appropriate $\varepsilon$ is chosen, such as $\varepsilon=10^{-5}$. Note that we omit the computation time and memory usage results, as this strategy does not impact these metrics.
For UCSD and DTDB dataset, we first compute the Reconstruction error of matrices clipped to different $\varepsilon$ values,
Figure 11: Average reconstruction error evaluated on different subspace dimensions of the UCSD dataset.
Figure 12: Average reconstruction error evaluated at each time step percentile of the DTDB dataset.
Figure 13: Average Reconstruction error over each dataset. Solid lines represent median values, and shaded areas indicate quartile values.
and show the results in Fig. 13 and Fig. 13, respectively. We can observe that the Reconstruction error of SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ are not distinguishable. However, if the $\varepsilon$ is chosen larger, i.e., $\varepsilon=10^{-2}$, the error becomes significantly larger, primarily because the learned system is too conservative and lacks expressivity.
Next, similar trend can be seen from the qualitative results of two selected sequences (DTDB dataset) shown in Fig. 16, where SC($\varepsilon=0$) and SC($\varepsilon=10^{-5}$) yield similar outcomes, while SC($\varepsilon=10^{-2}$) lacks expressivity, resulting in turning black. In addition, the quantitative results reported in Table. 4 also support this finding.
For DexManip dataset, the performance of SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ are also nearly distinguishable, as shown in Fig. 17 and Fig. 18 (Reconstruction error) and Table. 5 and Table. 6 (Success & Safety rates). Also, we can observe that $SC(\varepsilon=10^{-2})$ lacks the expressivity needed to generate diverse motions, leading to a significant low task success rate, although it maintains a 100% safety rate as all states converge to zero.
For Franka Panda dataset, though SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ still perform similarly, we observe that $SC(\varepsilon=10^{-2})$ also performs similarly, and even performs the best when using the fewest data points for training (see Fig. 19). We conjecture that this is due to the control inputs and feedback controller. In addition, all these three performs comparably in the long-term prediction (Fig. 20).
Figure 14: DTDB Sequence 185: A flushing toilet.
Figure 15: DTDB Sequence 205: A running stream.
Figure 16: Qualitative results of 2 sequences from DTDB dataset (same sequences as used in Fig. 6). Matrices are clipped to 1, clipped to $1-\varepsilon$, or not clipped.
Turning black ($\downarrow$) Blowing up ($\downarrow$) Normal ($\uparrow$) Moving ($\uparrow$)
LS 6.67% 45.96% 47.37% 62.11%
SC($\varepsilon=0$) 14.39% 0% 85.61% 54.39%
SC($\varepsilon=10^{-5}$) 14.74% 0% 85.26% 54.04%
SC($\varepsilon=10^{-2}$) 100.$0$% 0%
Table 4: Quantitative ratios of the characterization.
Tool Tool_short Reloc Reloc_with_failures
LS 55.5% 3.3% 96.0% 9.6%
SC($\varepsilon=0$) 94.0% 90.7% 93.8% 94.4%
SC($\varepsilon=10^{-5}$) 94.0% 89.0% 94.4% 93.8%
SC($\varepsilon=10^{-2}$) 24.7% 0% 66.7% 42.9%
Table 5: Task Success Rate for learned policy with varying $\epsilon$ values.
Tool Tool_short Reloc Reloc_with_failures
LS 0% 0% 100.0% 0.0%
SC($\varepsilon=0$) 100.0% 99.5% 100.0% 100.0%
SC($\varepsilon=10^{-5}$) 100.0% 99.5% 100.0% 100.0%
SC($\varepsilon=10^{-2}$) 100.0% 100.0% 100.0% 100.0%
Table 6: Safety Rate for the learned policy with varying $\epsilon$ values.
Figure 17: Average reconstruction error on the Tool and Reloc datasets. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
Figure 18: Average reconstruction error on the Tool_short and Reloc_with_failures datasets datasets. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
Figure 19: Rollouts of the end effector trajectory.
Figure 20: Distance to correct end effector position calculated at each time step.