### v4

# On the Surprising Effectiveness of Spectral Clipping in Learning Stable Linear Dynamics

Abstract
When learning stable linear dynamical systems from data, three important properties are desirable: i) predictive accuracy, ii) provable stability, and iii) computational efficiency. Unconstrained minimization of reconstruction errors leads to high accuracy and efficiency but cannot guarantee stability. Existing methods to enforce stability often preserve accuracy, but do so only at the cost of increased computation. In this work, we investigate if a straightforward approach can simultaneously offer all three desiderata of learning stable linear systems. Specifically, we consider a post-hoc approach that manipulates the spectrum of the learned system matrix that was computed using unconstrained least squares. We call this approach spectrum clipping (SC) as it involves eigen decomposition and subsequent reconstruction of the system matrix after clipping any eigenvalues that are larger than one to one (without altering the eigenvectors). Through comprehensive experiments involving two different applications and publicly available benchmark datasets, we show that this simple technique can efficiently learn highly-accurate linear systems that are provably-stable. Notably, we find that SC can match or outperform strong baselines while being orders-of-magnitude faster. We also show that SC can be readily combined with Koopman operators to learn stable nonlinear dynamics, such as those underlying complex dexterous manipulation skills involving multi-fingered robotic hands. Finally, we find that SC can learn stable robot policies even when the training data includes unsuccessful or truncated demonstrations.
Keywords: Linear Dynamical Systems, Stability, System Identification
1 Introduction
In spite of the growing dominance of deep neural networks, modeling and learning of linear dynamical systems (LDS) remain relevant due to the availability of analytical solutions, significantly lower computational burden, and ease of analyses and inspection. LDS continue to find use in diverse applications such as computer vision [1, 2, 3, 4] and time series prediction [5, 6]. Even when the underlying dynamics are nonlinear, researchers are increasingly utilizing tools from linear systems theory by leveraging the Koopman Operator theory [7, 8, 9] to approximately represent nonlinear systems as linear systems in a higher-dimensional lifted state space. These techniques have further increased the relevance and importance of learning LDS in complex applications, such as legged robotics [10, 11], robotic manipulation [12, 13, 14, 15], atmospheric sciences [16, 17, 18], and fluid dynamics [19, 20].
When time-series data associated with system evolution is available, learning LDS can be considered as a form of self-supervised learning and can be solved efficiently by minimizing a regression objective [21]. However, such unconstrained optimization methods tend to ignore important properties of the underlying systems, such as stability. Consequently, even if the underlying system is known to be stable, the stability of learned system could be highly sensitive to the training data distribution, potentially leading to unstable solutions and catastrophic long-term predictions.
A number of methods have been developed to enforce stability of learned LDS and reduce sensitivity to training data [22, 23, 2, 24, 25, 26, 27]. These methods take one of three approaches: a) minimizing the regression objective under stability constraints [22, 23], b) iterating between solving an unconstrained problem and refining the solution under stability constraints [2, 24], and c) iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective [25, 26, 27]. Regardless of the underlying approach, these methods inevitably introduce a considerable computational burden, potentially countering one of the primary advantages of leveraging linear techniques.
In this work, we investigate a simple yet surprisingly effective method, Spectrum Clipping (SC), to learn stable LDS from data while introducing negligible additional computations. SC involves a straightforward implementation: 1) compute the system matrix via unconstrained least-squares, 2) perform eigen decomposition of the system matrix and extract all eigenvalues, 3) clip eigenvalues with magnitudes larger than one to one, and 4) reconstruct the system evolution matrix using the clipped eigenvalues and the unchanged eigenvectors. Our investigation is inspired by prior works that have revealed the surprising effectiveness of linear techniques: controlling the singular values of linear layers improves training stability and generalization error [28, 29, 30]; retraining only the last linear layer improves robustness to spurious correlations [31]; and Low-Rank Adaptation (LoRA) dramatically improves efficiency [32].
SC enforces stability by construction without requiring constrained optimization or iterative updates. In contrast to existing approaches discussed above, SC can efficiently scale to high-dimensional systems (e.g., image prediction). Further, seminal works on the Koopman operator [33, 19, 9] reveal fundamental connections between the spectrum of the Koopman matrix and desirable properties in learned nonlinear systems. SC can be trivially combined with Koopman operators to enhance the predictive accuracy and stability of learned nonlinear dynamical systems (e.g., robotic systems).
We conducted comprehensive experiments involving benchmark datasets across different domains (two video prediction datasets [1, 34] and one robot model learning dataset [35]) and strong baselines (CG [2], WLS [24], and SOC [26]). Our results reveal that SC consistently achieves the lowest reconstruction error, and generates stable and more accurate long-term predictions. Notably, SC is able to achieve this level of performance while being orders-of-magnitude faster.
Additional experiments show that SC can be readily combined with Koopman operators to learn stable nonlinear systems (such as those that underlie complex dexterous manipulation skills for a multi-fingered robotic hand [13]) in mere seconds. Importantly, we show that SC enables the learned system to remain robustly stable and effective even the training data includes unsuccessful demonstrations or when the task horizon is truncated (characteristics are commonly observed when robot learning data is collected from humans [36, 37, 38]). In contrast, we find that unconstrained learning produces unstable LDS that lead to erratic and ineffective robot behaviors when such challenging characteristics are present in the training data.
In summary, our core contributions include: i) a simple post-hoc approach to learn stable linear dynamical systems, ii) evidence for its surprising ability to simultaneously offer reliable long-term predictions, high computational efficiency, and stability guarantees, iii) the ability to trade off between expressivity and stability, iv) an extension leveraging Koopman operators to accommodate stable nonlinear systems, and v) robustness against unsuccessful and truncated demonstrations.
2 Problem Formulation
Consider a stable discrete-time autonomous linear dynamical system $\mathrm{x}_{t+1}=A\mathrm{x}_{t}$, where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the system state at time $t$, and $A\in\mathcal{S}_{A}$ is the unknown system matrix, and $\mathcal{S}_{A}\subset\mathbb{R}^{n\times n}$ is the subspace that contains stable matrices: $\mathcal{S}_{A}\triangleq\{A\in\mathbb{R}^{n\times n}|\max\{|\lambda_{i}(A)|\}^{n}_{i=1}<=1\}$, where $\{\lambda_{i}(A)\}^{n}_{i=1}$ are the eigenvalues of the matrix $A$, and $|\lambda|$ is the $|\lambda|\coloneqq\sqrt{\text{Re}(\lambda)^{2}+\text{Im}(\lambda)^{2}}$.
Unconstrained Learning: Suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T}]$ containing the time evolutions of the above stable LDS. Defining $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, allows one to express the problem of learning the above LDS from a dataset $D$ as the following solving standard least-square regression problem [39, 40]:
$\hat{A}_{LS}=\arg\min_{\Theta}||Y-\Theta X||_{F}^{2},$ (1)
where $\Theta\in\mathbb{R}^{n\times n}$ are the learnable parameters and $||\cdot||_{F}$ is the matrix Frobenius norm. However, optimizing the Eq. (1) ignores the constraints on the matrix $A$, such as the system stability that is crucial for the accurate long-term prediction.
Learning Stable LDS: Our goal is to ensure that the learned system respects the stability constraints of the underlying unknown LDS, despite learning from the limited dataset $D$. To this end, we can rewrite the optimization problem Eq. (1) to include stability constraints:
$\hat{A}_{S}=\arg\min_{\Theta\in\mathcal{S}_{A}}||Y-\Theta X||_{F}^{2},$ (2)
While solving the optimization problem in Eq. (2) can yield the constrained optimal solution which is guaranteed to be stable, it is known to be highly non-convex and requires approximations [41, 26].
3 Spectrum Clipping: Enforcing Stability of Learned LDS
We evaluate a simple method, which we refer to as Spectrum Clipping (SC), to ensure the stability of learned LDS without having to directly solve the constrained optimization in Eq. (2).
Eigen Decomposition of Unconstrained Solution: First, we compute the system evolution matrix $\hat{A}_{LS}$ by optimizing the least square regression (Eq. (1)). Second, we perform eigen decomposition of the least squares solution: $\hat{A}_{LS}=M\Lambda M^{-1}$, where $\Lambda$ is a diagonal matrix with all the eigenvalues of $\hat{A}_{LS}$ on its diagonal, and $M$ is the modal matrix with the corresponding eigenvectors.
Spectral Clipping: Next, we identify all unstable eigenvalues in $\Lambda$ and clip their magnitudes to $1$: $\bar{\Lambda}_{i}=1,\forall i=j,\text{such that}|\Lambda_{j}|>1$ , and leave all other eigenvalues in $\Lambda$ unaltered: $\bar{\Lambda}_{i}=\Lambda_{i},\forall i=j,\text{such that}|\Lambda_{j}|\leq 1$. Finally, we reconstruct the system matrix based on the new $\bar{\Lambda}$ and the unaltered modal matrix $M$ to produce SC’s estimate of the system matrix: $\hat{A}_{SC}\triangleq M\bar{\Lambda}M^{-1}$.
Note that the above post-hoc clipping procedure enforces stability constraints by ensuring that $\hat{A}_{SC}\in\mathcal{S}_{A}$ without requiring constrained optimization. Indeed, it is natural to suspect that this seemingly-naive approach would result in significantly poorer, albeit stable, approximations and worse prediction accuracies. However, as our experiments demonstrate, SC surprisingly leads to similar, if not better, accuracy compared to existing constrained optimization techniques.
Ensuring Lyapunov Stability: Enforcing Since SC clips eigenvalues to 1, the resulting system may not be strictly stable in the sense of Lyapunov. If desired, SC can be trivially extended to ensure Lyapunov stability by merely clipping the magnitude of the eigenvalues to marginally less than $1$. We formalize this by clipping the magnitudes to $1-\varepsilon$ (instead of $1$), where $\varepsilon\in[0,1]$ and Lyapunov stability is ensured when $\varepsilon>0$.
Expressivity-Stability Tradeoff: We note that the choice of $\varepsilon$ above represents an important tradeoff between the system’s expressivity and stability. In particular, choosing $\varepsilon=0$ will yield marginally-stable but highly-expressive systems. In contrast, choosing $\varepsilon>0$ will ensure strictly-stable system, with increasing values of $\varepsilon$ yielding faster exponential convergence and diminishing expressivity (see Appendix. E for experimental validation).
Dealing with Non-diagonalizable Unconstrained Solutions: A potential concern about the above approach is that $\hat{A}_{LS}$ could be non-diagonalizable. However, it is well known that one can always first find a diagonalizable matrix arbitrarily close to the original non-diagonalizable matrix (see Lemma. 3.1).
Lemma 3.1.
For any $\gamma>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. $||\hat{A}_{LS}-\tilde{A}_{LS}||_{1}<\gamma$.
Proof.
Lemma 3.1 can be proved by the fact that the set of real matrices that are diagonalizable over $\mathbb{C}$ is dense in $\mathbb{R}^{n\times n}$. The details can be found from the matrix perturbation theory [42, 43]. ∎
The above result can be trivially extended to show that the evolution of the system with a non-diagonalizable matrix and that of the system with the arbitrarily-close diagonalizable matrix can be made arbitrarily close (see proof in Appendix. A). As such, the SC method can be applied to the arbitrarily-close diagonalizable matrix $\tilde{A}_{LS}$ to enforce stability of systems with non-diagonalizable matrices without significant approximation errors.
4 Empirical Analyses
In this section, we evaluate SC in terms of computation efficiency, stability, and expressivity when learning LDS from data and compare its performance against existing methods. Note that for all experiments, we clip the spectrum to $1$ (see Appendix. E for the effects of different $\varepsilon$ values).
4.1 Experimental Design
Computation Platform:
We conducted all experiments on Alienware Aurora R13, with 10-core 12th Gen Intel Core i7-12700KF 1.0-GHz CPU and a 32G RAM.
Baselines:
i) LS: vanilla least squares, ii) (CG): constraint generation [2], iii) WLS: weighted least squares [24], and iv) SOC: a characterization approach [26, 27].
Datasets:
We evaluate all approaches on the following datasets.
UCSD [1]: A dataset of 254 highway traffic videos. Each video has 48-52 frames and the size of each frame is $48\times 48$. To lower the dimension, we apply the SVD method on each sequence to obtain a subspace dimensions $r\in\{3,30\}$, as similarly done in [26].
DTDB [34]: A dataset of 285 dynamic texture movements, such as turbulence, wavy motions, etc. Each video consists of varying frames, and we also apply the SVD method to lower the state dimension to $r=300$.
Metrics:
We evaluate all approaches in terms of: i) Computation time: Time taken to train a stable LDS, ii) Memory Usage: Memory used during the entire computation process, and iii) Reconstruction error: The $L1$ distance between generated state sequences and ground truth states.
4.2 SC significantly reduces computation & memory burden
In this subsection, we compare the training efficiency among each approach. We first report the average computation time across the 254 input sequences of the UCSD dataset at each dimension in Fig. 1 (a). The results show that SC is significantly more time-efficient than WLS, CG, and SOC. This is because that SC only needs the eigen-decomposition and eigen-recomposition to analytically compute the stable LDSs, while others require to iteratively find a solution. Note that though LS is the most time-efficient, it can not provide with stability.
In Fig. 1 (b), we then report the results on the DTDB dataset. The results also indicate that the computation time of SC is significantly and consistently shorter than that of SOC, showing that SC is more scalable w.r.t. the input dimensions. Similarly, the memory usage results reported in Fig. 1 (c) and Fig. 1 (d) suggest that SC requires significantly less memory compared to CG, WLS, and SOC.
Figure 1: (a): Computation time on UCSD dataset (solid lines indicate the median); (b): The boxplot of the average computation time on DTDB dataset; (c): Memory usage on UCSD dataset (solid lines indicate the median); (d): The boxplot of the average memory usage on DTDB dataset.
LS CG [2] WLS [24] SOC [26, 27] SC
Time complexity -- $O(n^{6})$ $O(n^{6})$ $O(n^{3})$ $O(n^{3})$
Space complexity $O(n^{2})$ $O(n^{4})$ $O(n^{4})$ $O(n^{2})$ $O(n^{2})$
Table 1: Space and Time complexity of each method. $n$ is the system dimension. For time complexity, since all methods modify the original LS solution to enforce stability, we compare only the cost incurred after solving LS. For iterative methods (CG, WLS, and SOC), the shown time complexity corresponds to a single iteration.
In addition to empirical results on the two datasets, we report the time and space complexity of each method in Table. 1, further supporting the observation that SC is as or more efficient in both time and memory than the other methods. For time complexity, since all methods modify the original LS solution to enforce stability, we compare only the cost incurred after solving LS. Note that CG, WLS, and SOC are iterative algorithms, and the table reports the time complexity per iteration. In contrast, SC is non-iterative, and its overall time complexity is shown.
4.3 SC effectively identifies the underlying LDSs
In this subsection, we compare the prediction accuracy of the identified LDSs by each approach. We first present the average reconstruction error across all states at each time step for the UCSD dataset in Fig. 2 (left). For illustrative purposes, we select $r=3,20$ and $30$ as examples. It is clearly shown that as the input dimension increases (e.g., $r=30$), LS fails to reconstruct the input states, with the error increasing significantly over time. A similar trend can be observed in Fig. 3, where we show the average reconstruction error over all states and time steps of each method at each training dimension. In both cases, SC consistently achieves results comparable to CG, WLS, and SOC, while being orders of magnitude faster for computation and highly efficient in memory usage (see Section. 4.2).
We then show the average reconstruction error for the DTDB dataset in Fig. 2 (right). Note that due to the high dimension of the state space in this dataset, only LS, SOC, and SC are applicable. Both CG and WLS involve solving a Kronecker product, which has a significantly large space complexity (i.e., $O(n^{4})$ shown in Table. 1, $n$ is the size of the state space), leading to memory overflow.
Given the varying time steps of the input sequences, the reconstruction error is averaged over the time percentiles. The results indicate that SC outperforms SOC since it achieves a lower reconstruction error. Note that though the LS approach achieves the lowest error within the training data horizon, it does not guarantee stability, which is crucial for long-term prediction (see Section. 4.4).
Figure 2: Left figure: average reconstruction error evaluated on different subspace dimensions of the UCSD dataset. Shaded areas indicate quartile values over all sequences. Right figure: average reconstruction error evaluated at each time step percentile of the DTDB dataset. Shaded areas indicate quartile values over all sequences.
4.4 SC enables stable long-term predictions
Figure 3: Average reconstruction error evaluated on different subspace dimensions of the UCSD dataset.
In this subsection, we demonstrate that the SC approach learns LDSs capable of making long-term dynamic predictions that extend beyond the training data horizon. To verify this, we extend the rollout horizon to $T=3000$ time steps for the 285 LDSs learned from the DTDB dataset. In this case, an accurate LDS should be able to repeat the dynamical pattern learned from the training data without either blowing up or converging during the extended rollout horizon.
We first select two example sequences (i.e., flushing toilet and running stream) and show the qualitative results in Fig. 6. These results indicate that LS approach mostly learns the LDSs that blow up during the extended horizon, as it does not provide with stability guarantee. Conversely, the SOC approach tends to learn LDSs that converge to zeros, indicating a lack of dynamic behavior. In Table. 2, we report the spectral radius of the learned system matrix for each method. The results support the conclusion that LS leads to blow-ups due to a spectral radius greater than one, while SOC results in vanishing values as its spectral radius remains strictly less than one.
LS SOC [26, 27] SC
DTDB Sequence 185 1.0025 0.9999 1.0000
DTDB Sequence 205 1.0022 0.9999 1.0000
Table 2: Spectrum radii of the system matrices learned with each method.
Figure 4: DTDB Sequence 185: A flushing toilet.
Figure 5: DTDB Sequence 205: A running stream.
Figure 6: Qualitative results for 2 example sequences from the DTDB dataset.
We then define several quantitative evaluation metrics, which are divided into two categories.
Quality (NIQE score [44]): A commonly used image quality evaluation metric. This metrics helps determine if the predictions of the LDSs are still meaningful. Note that images with extremely low quality (for example, when most pixel values converge to 0 or blow up) can produce unbounded NIQE scores. We substitute those unbounded scores with $10^{5}$ to provide a bounded statistical summaries.
Dynamics (Moving Ratio  [^1] Moving is defined as $||F_{M}-F_{M-2}||_{1}>T$, where $\{F_{i}\}_{i=1}^{M}$ is a sequence of $M$ image frames and $T=9$.): The ratio of sequences in which the final frames still exhibit the dynamic pattern This characterization aims to determine if the system evolution of the LDSs still reflects the dynamic pattern.
Average NIQE ($F=500$) ($\downarrow$) Average NIQE ($F=3000$) ($\downarrow$) Moving ($\uparrow$)
LS 904.6265 6355.5817 62.11%
SOC [26, 27] 832.8063 4243.4279 22.11%
SC 18.8769 1075.0140 54.39%
[…]
Table 3: Quantitative metrics evaluating the predictions of each method on the DTDB dataset.
In Table. 3, we report the quantitative ratios in terms of the defined metrics over the 285 sequences in the DTDB dataset. Note that the values for LS method are marked in gray as this method does not guarantee stability. The results indicate that SC best preserves long-term dynamic patterns while maintaining high prediction accuracy, as evidenced by its superior image quality—particularly at 500-step predictions, which are nearly comparable to the training set images.
4.5 SC’s benefits extend to nonlinear systems when combined with the Koopman Operator
In this subsection, we demonstrate that when combined with the Koopman Operator theory [7], SC can also be effectively and efficiently applied to highly nonlinear robot manipulation tasks. For this, we introduce a highly nonlinear robot manipulation dataset.
DexManip [13, 45]: A dataset of Adroit Hand performing two dexterous manipulation tasks in MUJOCO simulation environment [46]. Though the original state evolution is nonlinear, but as shown in [13], via lifted to a higher-dimension Hilbert space (e.g., $r\approx 750$), the system can be considered linear in the lifted state space. Two tasks are briefly described as follows: i) Tool Use: The robot hand is tasked with picking up the hammer and driving the nail into the board placed at a random height, and ii) Object Relocation: The robot hand is tasked with moving a cylinder to a random target location (see Appendix. B for the details of each dataset). Similar to the DTDB dataset, only LS, SOC, and SC are applicable due to the high dimensionality of the lifting space.
Computation Time (sec) Memory Usage (MB) Task Success Rate (%) Safety Rate (%)
Datasets LS SOC SC LS SOC SC LS SOC SC LS SOC SC
Tool 0.6 192.8 0.8 224.6 331.7 257.2 55.5 0.0 94.0 0.0 100.0 100.0
Reloc 0.7 6608.1 0.8 215.3 321.3 247.5 96.0 48.6 93.8 100.0 100.0 100.0
Table 4: SC not only alleviates the computational and memory burden on the DexManip datasets, but also enables successful and safe robot manipulation.
We first report the computation time and memory usage in Table. 4. As seen, SC also alleviates the computational and memory burden compared to the SOC method. Note that the values for LS method are marked in gray as this method does not guarantee stability. We then show the average reconstruction error on the DexManip datasets in Fig. 7 (a) & (b), where SC demonstrates the best performance and LS results in significantly larger errors on the Tool dataset due to its lack of stability.
Figure 7: (a): Average reconstruction error on Tool dataset; (b): Average reconstruction error on Reloc dataset; (c): Average reconstruction error on Tool_short dataset; (d): Average reconstruction error on Reloc_with_failures dataset. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
To further underscore the importance of ensuring stability when learning LDSs for robotic dexterous manipulation tasks in the DexManip datasets, we also compute the task success rates and the safety rates (both reported in Table. 4) for each task when tracking the generated robot hand trajectories. The definition of these two metrics are included in Appendix. B.
For both tasks, SC learns stable LDSs that achieve the highest task success rates and complete the tasks safely. In contrast, SOC underperforms due to the fact it learns overly conservative LDSs that lack the ability to generate diverse trajectories. Note that while LS seems to achieve the desired performance on the Reloc task, its effectiveness drops significantly when trained on a corrupted dataset as shown in Section. 4.6.
4.6 SC can learn stable LDSs from shorter and corrupted datasets
In real-world settings, the collected demonstrations may be imperfect, unlike those in the original DexManip datasets. Hence, we introduce two additional datasets: (i) Tool_short, where each demonstration is shorter, posing challenges in learning stable LDSs with a limited horizon, and (ii) Reloc_with_failures, which includes a few failed demonstrations (see Appendix. B for more details).
Computation Time (sec) Memory Usage (MB) Task Success Rate (%) Safety Rate (%)
Datasets LS SOC SC LS SOC SC LS SOC SC LS SOC SC
Tool_short 0.3 3470.2 0.4 80.2 280.5 112.8 3.3 0.0 90.7 0.0 100.0 100.0
Reloc_with_failures 0.6 4564.6 0.9 216.1 322.1 248.4 9.6 0.0 94.4 0.0 100.0 100.0
Table 5: SC can efficiently learn stable manipulation skills from shorter and corrupted datasets.
In Table. 5, we report the computational time and memory usage, and it is evident that SC is more efficient than SOC. In Fig. 7 (c) & (d), we show the average reconstruction error over the original uncorrupted datasets. From these results, we can observe that the LDSs learned by the LS approach tend to be highly unstable, as evidenced by the significantly increasing error over time. This indicates the importance of enforcing stability in learned LDSs to effectively handle corrupted demonstrations.
A similar conclusion can be drawn from the task success rates and safety rates (both reported in Table. 5), where SC demonstrates significantly superior performance.
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in [33], we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset.
Specifically, for each eigenvalue $\lambda_{i}\ (i=1,...,n)$ of Koopman matrix $K$, we can find corresponding eigenvector $\mathbf{v_{i}}$ (i.e., the mode) and eigenvector $\mathbf{w_{i}}$ of adjoint matrix $K^{*}$, such that $K\mathbf{v_{i}}=\lambda_{i}\mathbf{v_{i}}\text{, and }K^{*}\mathbf{w_{i}}=\overline{\lambda_{i}}\mathbf{w_{i}}$. Then following [33], we construct the corresponding eigenfunction $\varphi_{i}:\mathbb{R}^{n}\rightarrow C$, defined as $\varphi_{i}(\mathbf{z})=\langle\mathbf{z},\mathbf{w}_{i}\rangle$. By doing so, the rollout of eigenfunctions can be represented by $\varphi_{i}(K\mathbf{z})=\lambda_{i}\varphi_{i}(\mathbf{z})$, and the states can be reconstructed by $\mathbf{z_{k}}=\sum_{i=1}^{n}\lambda_{i}^{k}\varphi_{i}(\mathbf{z_{0}})\mathbf{v_{i}}$. Based on this formula, specific modes can be selected for rollouts.
In our experiments, we first computed the Koopman matrix $K_{f}$ from $Reloc\_with\_failures$ dataset. We then sorted the eigenvalues by their norms from largest to smallest and divided them into two subsets for rollouts: (i) Unstable Set: the first 52 Koopman modes with unstable eigenvalues, which have norms greater than 1, and (ii) Stable Set: the remaining 707 modes. We then visualized the robot hand motions of each subset via the eigenfunction rollouts in Fig. 8, respectively (see the supplementary video for full policy rollouts). It can be seen that the rollout of the Unstable Set before clipping clearly generate unstable hand motions (see upper line of the green box), which results in the low success rate and safety rate when combined with Stable Set for the final task performance (as shown in Table. 5). Further, we clipped the matrix $K_{f}$ and visualized the rollouts of the clipped Unstable Set in Fig. 8. It is obvious that after clipping, the robot hand motions become stable (see bottom line of the green box), enabling the composite system to safely complete the task (as also shown in Table 5). The full policy rollout is also available in the supplementary video.
Figure 8: Visualizations of rollout with two subsets of the Koopman modes. The green box represents rollouts using Unstable Set (norm of eigenvalue larger than 1). The blue box represents rollout with the rest Stable Set. In the green box, the upper line shows the rollout before clipping and the lower line shows it after clipping. It is obvious that after clipping, the rollout becomes stable, thereby allowing the composite rollout of both sets to safely complete the task (as shown in Table 5. The full policy rollout is available in the supplementary video.
4.7 SC can identify systems with control inputs
In this subsection, we demonstrate the comparable capability of SC in identifying stable LDSs with control inputs (see Appendix. C for details). To this end, we first introduce another dataset that includes the control commands.
Franka Panda [35]: A dataset of Franka Panda Robot Arm’s movements. The robot state space has 17 dimensions ($r=17$), including the coordinates of end effector ($\mathbb{R}^{3}$), joint angle values ($\mathbb{R}^{7}$), and joint angular velocities ($\mathbb{R}^{7}$), and the control input has $\mathbb{R}^{7}$ joint torques. This dataset consists of 3100 time steps of control inputs (randomly generated) and the corresponding robot states.
To examine how dataset size affects performance, we adjust the dataset size to $d=100,2000$ and $3000$ samples. For each size, we randomly select five subsets from the whole dataset, resulting in totally 15 subsets. We then train LDSs, i.e., the system matrix $A$ and the control matrix $B$, using different approaches on all 15 subsets.
We first evaluate the long-term prediction accuracy of each approach. For this, we first command the random control inputs in simulation with ground-truth system dynamics to generate test trajectories. We then rollout the trained $A$ and $B$ matrices with the same control inputs, and then compute the average $L2$ distance between the predicted and ground-truth end-effector positions for each method. From the results shown in Table. 6, We observe that while SC performs worse than CG and SOC when $d=100$, its performance becomes comparable as $d$ increases. Notably, SC achieves this with significantly lower computation time and memory usage.
Additionally, we use the trained $A$ and $B$ matrices to construct an LQR controller, and use such controller to track a figure-8 pattern in simulation (similar in [26]). How accurately the figure is tracked indicates how well the learned LDS aligns with the actual dynamical system with control inputs. For each approach, we repeat this process five times with different initial positions of the end effector. We show the qualitative tracking performances in Appendix. D, where SC demonstrates performance comparable to that of other methods.
LS CG [2] WLS [24] SOC [26, 27] SC
$d=100$ $0.0748\pm 0.1218$ $0.0166\pm 0.0066$ $0.3390\pm 0.2059$ $0.0225\pm 0.0131$ $0.0693\pm 0.1370$
$d=2000$ $0.0207\pm 0.0173$ $0.0154\pm 0.0058$ $0.3051\pm 0.1853$ $0.0154\pm 0.0058$ $0.0183\pm 0.0138$
$d=3000$ $0.0164\pm 0.0083$ $0.0155\pm 0.0058$ $0.2845\pm 0.1501$ $0.0155\pm 0.0058$ $0.0156\pm 0.0057$
Table 6: Average prediction error of the end-effector’s position (unit: $m$).
5 Limitations and Future Work
Though spectrum clipping (SC) demonstrates its efficacy on identifying the stable LDSs over various datasets, there are still opportunities for future study. First, the performance on the dataset with control inputs is not yet fully comparable to other state-of-the-art approaches, indicating potential for algorithmic improvements. Second, while we empirically demonstrate the effectiveness of SC, a deeper theoretical analysis will likely offer valuable insights.
6 Conclusion
In this paper, we demonstrate that Spectrum Clipping is an efficient data-driven method for identifying stable LDSs. Through extensive experiments across multiple datasets, we show that it is not only significantly more time- and memory-efficient but also achieve superior performance. Furthermore, we demonstrate that it also excels in long-term dynamics prediction. Additionally, it can be applied to the highly nonlinear robot manipulation tasks with the Koopman Operator theory and the linear dynamical systems with control inputs.
Appendices
Appendix A Proof for Arbitrarily Small System Evolution Error
Theorem A.1.
For any $\eta>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, we can always find its diagonalizable counterpart $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$.
Proof.
$\tilde{A}_{LS}$ can be denoted as
$\tilde{A}_{LS}=\hat{A}_{LS}+E$ (3)
where $E\in\mathbb{R}^{n\times n}$ is the residual matrix and, from Lemma. 3.1, $||E||_{1}$ can be arbitrarily small (i.e., $||E||_{1}<\gamma$). Then we have
$\tilde{A}_{LS}^{k}\mathbf{z}=(\hat{A}_{LS}+E)^{k}\mathbf{z}=\hat{A}_{LS}^{k}\mathbf{z}+\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}$ (4)
Therefore,
$\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}=-\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}$ (5)
Using triangle inequality, we have
$||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}=||\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}||_{1}<=\sum_{i=1}^{k}||{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}||_{1}$ (6)
Since $\hat{A}_{LS}$, $k$ and $\mathbf{z}$ are all bounded, there exists a bounded $m>0$, s.t.
$||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<m\gamma$ (7)
Let $\eta=m\gamma$. Note that since Lemma. 3.1 guarantees the existence of a diagonalizable matrix $\tilde{A}_{LS}$ that satisfies an arbitrarily small $\gamma$, it follows that $\eta=m\gamma$ can also be made arbitrarily small. Therefore, for any $\eta>0$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$. ∎
Appendix B Details of DexManip tasks
DexManip is a dataset collected when using Adroit Hand to conduct two different dexterous manipulation tasks in MUJOCO simulation environment. In this dataset, the states of hand and states of object is captured, concatenated and lifted to a higher-dimension Hilbert space at each time step (see [13]), so that the hand-object system becomes an LDS in the Hilbert space. Thereafter, we use a simple PD controller to track the generated hand trajectories in the simulator. The tasks are detailed as follows.
Tool Use: This task is mainly about picking up the hammer to drive the nail into the board placed at some height. The hand and object states are lifted to a higher-dimensional $z$ states and different approaches are used to learn the LDSs in the $z$ state space. There are two types of datasets:
  • Tool: This dataset consists of 200 rollouts that continues after the nail is driven into the board, so that the hammer stays stable in the air ($T=100$). This is the original setting proposed in [47].
  • Tool_short: This dataset consists of 200 rollouts that end right after the nail is driven in ($T=35$).
This task is considered i) successful if at last time step (note that during evaluation, the policy will always run for 100 time steps, similar to the Tool dataset), the Euclidean distance between the final nail position and the goal nail position is smaller than 0.01m [13], and ii) safe if the final distance from the hammer to the goal nail position is smaller than 0.3m, indicating that the robot hand neither drifts away nor throws away the hammer.
Object Relocation: This task is mainly about moving a cylinder to a randomized target location. Similarly, all approaches are operated in the higher-dimensional $z$ state space. The two types of datasets are:
  • Reloc: This dataset consists of 177 rollouts that objects are all moved to the target position successfully in the end ($T=100$).
  • Reloc_with_failures: In addition to the Reloc dataset, this one also includes the failure cases where the objects are failed to move to the target position in the end ($T=100$).
At each time step, if the Euclidean distance between the current cylinder position and the target position is smaller than 0.10m, then we have $\rho(t)=1$. This task is considered i) successful if $\sum_{t=1}^{T=100}\rho(t)>10$ (this criterion is the same as that used in [13]), and ii) safe if the robot hand is still near the target position (i.e., the distance is smaller than 0.5m) at final time step, indicating that the robot hand does not drifts away.
Appendix C Spectrum Clipping for Controlled System
Consider a discrete-time controlled linear system as
$\mathrm{x}_{t+1}=A\mathrm{x}_{t}+B\mathrm{u}_{t},$ (8)
where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n},\mathrm{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ are the system state and control signal at time $t$, respectively. Similarly, $A\in\mathcal{S}_{A}$ is the unknown system matrix, and $\mathcal{S}_{A}\subset\mathbb{R}^{n\times n}$ is the subspace that contains stable matrices: $\mathcal{S}_{A}\triangleq\{A\in\mathbb{R}^{n\times n}|\max\{|\lambda_{i}(A)|\}^{n}_{i=1}<=1\}$, and $B\in\mathcal{S}_{B}$ is the unknown control matrix, and $\mathcal{S}_{B}\subset\mathbb{R}^{n\times m}$.
Unconstrained Learning: Now, suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{u}_{1},\mathrm{x}_{2},\mathrm{u}_{2},\cdots,\mathrm{x}_{T}]$. Let $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, and $U=[\mathrm{u}_{1},\mathrm{u}_{2},\cdots,\mathrm{u}_{T-1}]\in\mathbb{R}^{m\times T-1}$, then to learn the linear system from the dataset $D$ is to minimize the similar regression objective:
$\hat{A}_{LS},\hat{B}=\arg\min_{\{\Theta,\Gamma\}}||Y-AX-BU||_{F}^{2},$ (9)
where $\Theta\in\mathbb{R}^{n\times n}$ and $\Gamma\in\mathcal{S}_{B}$ are both the learnable parameters, and $||\cdot||_{F}$ is the matrix Frobenius norm. Similarly, optimizing Eq. (9) does not guarantee stability.
Learning Stable LDS with Control Inputs: To this end, we can rewrite the optimization problem Eq. (9) to include stability constraints:
$\hat{A}_{S},\hat{B}=\arg\min_{\{\Theta\in\mathcal{S}_{A},\Gamma\}}||Y-AX-BU||_{F}^{2},$ (10)
To find the approximate solution to Eq. (10), we first compute $\hat{A}_{LS}$ and $\hat{B}$ by optimizing the least square regression (Eq. (9)). Then we apply SC method to $\hat{A}_{LS}$ and obtain $\hat{A}_{SC}$ as described in Section. 3, while keeping $\hat{B}$ unchanged.
Appendix D Qualitative performance on the Franka Panda dataset
From the results shown in Fig. 9, we can observe that CG and SOC overall exhibit the best performance, especially when using only 100 training samples. However, as dataset size increases, the performance of SC is significantly improved, and become comparable to both CG and SOC, while still being orders of magnitude faster for training.
Figure 9: Rollouts of the end effector trajectory. Each subplot contains 25 trajectories, representing 5 different initial positions with matrices trained on 5 subsets.
Appendix E Tradeoff between stability and expressivity
As mentioned in Section. 3, we can also select the $\varepsilon$ values to ensure that the resulting linear system meets the criteria for well-known Lyapunov stability. Through the experiments conducted on UCSD, DTDB, DexManip, and Franka Panda datasets, we demonstrate that this strategy does not only guarantee the Lyapunov stability, but also achieve comparable performance, provided that an appropriate $\varepsilon$ is chosen, such as $\varepsilon=10^{-5}$. Note that we omit the computation time and memory usage results, as this strategy does not impact these metrics.
For UCSD and DTDB dataset, we first compute the Reconstruction error of matrices clipped to different $\varepsilon$ values,
Figure 10: Left figure: average reconstruction error evaluated on different subspace dimensions of the UCSD dataset. Solid lines represent median values, and shaded areas indicate quartile values. Right figure: Average reconstruction error evaluated at each time step percentile of the DTDB dataset. Solid lines represent median values, and shaded areas indicate quartile values.
Figure 11: Average reconstruction error evaluated on different subspace dimensions of the UCSD dataset.
and show the results in Fig. 10. We can observe that the Reconstruction error of SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ are not distinguishable. However, if the $\varepsilon$ is chosen larger, i.e., $\varepsilon=10^{-2}$, the error becomes significantly larger, primarily because the learned system is too conservative and lacks expressivity.
Next, similar trend can be seen from the qualitative results of two selected sequences (DTDB dataset) shown in Fig. 14, where SC($\varepsilon=0$) and SC($\varepsilon=10^{-5}$) yield similar outcomes, while SC($\varepsilon=10^{-2}$) lacks expressivity, resulting in turning black. In addition, the quantitative results reported in Table. 7 also support this finding.
For DexManip dataset, the performance of SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ are also nearly distinguishable, as shown in Fig. 15 and Fig. 16 (Reconstruction error) and Table. 8 and Table. 9 (Success & Safety rates). Also, we can observe that $SC(\varepsilon=10^{-2})$ lacks the expressivity needed to generate diverse motions, leading to a significant low task success rate, although it maintains a 100% safety rate as all states converge to zero.
For Franka Panda dataset, though SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ still perform similarly, we observe that $SC(\varepsilon=10^{-2})$ also performs similarly, and even performs the best when using the fewest data points for training (see Fig. 17). We conjecture that this is due to the control inputs and feedback controller. In addition, all these three performs comparably in the long-term prediction (Table. 10).
Figure 12: DTDB Sequence 185: A flushing toilet.
Figure 13: DTDB Sequence 205: A running stream.
Figure 14: Qualitative results of 2 sequences from DTDB dataset (same sequences as used in Fig. 6). Matrices are clipped to 1, clipped to $1-\varepsilon$, or not clipped.
Average NIQE ($F=500$) ($\downarrow$) Average NIQE ($F=3000$) ($\downarrow$) Moving ($\uparrow$)
LS 904.6265 6355.5817 62.11%
SC($\varepsilon=0$) 18.8769 1075.0140 54.39%
SC($\varepsilon=10^{-5}$) 18.8769 1075.0139 54.04%
SC($\varepsilon=10^{-2}$) 706.0687 inf 0%
[…]
Table 7: Quantitative metrics evaluating the predictions with varying $\varepsilon$ values on the DTDB dataset.
Tool Tool_short Reloc Reloc_with_failures
LS 55.5% 3.3% 96.0% 9.6%
SC($\varepsilon=0$) 94.0% 90.7% 93.8% 94.4%
SC($\varepsilon=10^{-5}$) 94.0% 89.0% 94.4% 93.8%
SC($\varepsilon=10^{-2}$) 24.7% 0% 66.7% 42.9%
Table 8: Task Success Rate for learned policy with varying $\epsilon$ values.
Tool Tool_short Reloc Reloc_with_failures
LS 0% 0% 100.0% 0.0%
SC($\varepsilon=0$) 100.0% 99.5% 100.0% 100.0%
SC($\varepsilon=10^{-5}$) 100.0% 99.5% 100.0% 100.0%
SC($\varepsilon=10^{-2}$) 100.0% 100.0% 100.0% 100.0%
Table 9: Safety Rate for the learned policy with varying $\varepsilon$ values.
Figure 15: Average reconstruction error on the Tool and Reloc datasets. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
Figure 16: Average reconstruction error on the Tool_short and Reloc_with_failures datasets datasets. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
Figure 17: Rollouts of the end effector trajectory.
LS SC($\varepsilon=0$) SC($\varepsilon=10^{-5}$) SC($\varepsilon=10^{-2}$)
$d=100$ $0.0748\pm 0.1218$ $0.0693\pm 0.1370$ $0.0686\pm 0.1344$ $0.0193\pm 0.0074$
$d=2000$ $0.0207\pm 0.0173$ $0.0183\pm 0.0138$ $0.0182\pm 0.0132$ $0.0175\pm 0.0062$
$d=3000$ $0.0164\pm 0.0083$ $0.0156\pm 0.0057$ $0.0156\pm 0.0057$ $0.0176\pm 0.0062$
Table 10: Average prediction error of the end-effector’s position (unit: $m$).